# Radial Gradients and Intrinsic Scatter in MaNGA Galaxies

### v1

@@ 3.2 Model
Naturally, the results of galactic spectral analysis depend on the stellar population model used. The stellar population model utilized for this work is based on Worthey et al. (2014) (hereafter Worthey model) with the foundational framework established by Worthey (1994). These models use stellar evolutionary isochrones combined with a stellar initial mass function (IMF) to estimate the distribution of stars within the log L versus log Teff diagram. Fluxes are assigned to each star bin, along with empirical estimates of the absorption feature indices.
@@
For any stellar population model three major ingredients are: stellar spectral library, set of isochrones, and initial mass function (IMF). The Worthey model is based on different stellar spectral libraries: 1. a variant of Worthey et al. (1994) Lick spectra 2. the MILES spectral library (Sánchez-Blázquez et al., 2006) 3. the Indo-US library by Valdes et al. (2004) 4. the IRTF spectral library (Rayner et al., 2003) 5. the ELODIE spectral library (Prugniel & Soubiran, 2001). An index fitting function was used to fit the stellar indices as a function of Teff, log($g$), and [Fe/H] and the functions are incorporated into the models. Marigo et al. (2008) isochrones were used with element sensitivity added to the isochrones according to Worthey et al. (2022). The results from the index fitting functions was associated with this isochrones adopting a Kroupa IMF (Kroupa, 2001). To generate an integrated-light index strength, the flux and continuum values for each ”star” along the isochrone are first calculated, with each star’s contribution weighted according to its number. These weighted values are then summed across the entire isochrone. After the summation, the resulting data is reformulated into a final index value, representing the combined (simple) stellar population’s spectral characteristics. An additional weighted sum is performed for this work. Our model also incorporates [Fe/H]-dependent abundance distribution via the abundance distribution function (ADF) (Tang et al., 2014; Worthey et al., 2005). This allows us to model metallicity-composite stellar populations, but still with a single burst age. In this work we adopt the ‘normal’ width ADF with FWHM = 0.65 dex from Tang et al. (2014).
With a grid of composite models in hand, an inversion program (compfit, to fit observations with a grid of composite models) was developed to determine the optimal set of abundance parameters for a specified set of indices. It is an improved version of the inversion code used in Worthey et al. (2014).
@@
Here are a few notes on how to interpret the inverted-model results. The final abundance for each element is listed here as the peak (mode) of the ADF, not any kind of SSP-equivalent abundance. Light elements (C, N, O, Si, Na) are assumed to be related to Fe and $Z$ in a single ratio for each solution (that is, no trend of [$\alpha$/Fe] with Fe or $Z$ was assumed). Furthermore, due to the subtlety of the spectra features affected by Si and O, inclusion of these elements gives noisy solutions. Therefore, these two elements were tied to Mg, which strongly affects the spectrum. That is, [Mg/Fe]=[Si/Fe]=[O/Fe] always. This matters because isochrone temperatures vary with O and Si abundance, and there is a strong O-age degeneracy (Worthey et al., 2022). Elements not mentioned, such as Ni, Cr, Pd, or U, were similarly tied in lockstep to Fe. In this paper, we refer to these unnamed elements by the symbol ‘R.’ Since, for this paper only, we did not allow Fe to vary on its own, R is equivalent to Fe. For example, [Mg/R] = [Mg/Fe]. Helium abundance has a non-adjustable trajectory with $Z$ determined by the isochrone set. This is out of pragmatic necessity, since He cannot be measured in the galaxies studied here, but should be mentioned since it puts an irreducible systematic error into the results.
@@ 4 Results
In this section we present the results on the age, [Fe/H], and C, N, Na, and Mg abundances and their radial gradients. For the scope of this paper, the single-burst-equivalent mean age, [Fe/H], [C/R], [N/R], [Na/R], and [Mg/R] are collectively referred to as stellar population parameters. We first present the stellar population parameters as a function of R/Re till R/Re=0.7 in §$\S$4.1. Average parameters binned in $\sigma$ are listed in Table 2 and parameters binned by [Fe/H] are presented in Table 3. For both of these tables, the values tabulated are the mean of the values from the 3rd annulus of galaxies within a specific bin. In §$\S$4.2 we present the result on radial gradients in the stellar population parameters.
@@ 4.1 Stellar Population Parameters
[…]
@@ 4.1.1 Age
Since mass is proportional to the square of the velocity dispersion but only linear with galaxy size (Bender et al., 1992), $\sigma$ is an imperfect but adequate proxy for the mass of the galaxy. Fig. 6 shows how the log of mean age (Gyr) changes with $\sigma$. The top panel shows mean age of each annulus (annuli are color coded) plotted against log of $\sigma$. The $\sigma$ values on the x-axis are the middle values of the each $\sigma$ bin (described in 2). The bottom panel shows the mean age for each of the galaxy at their 3rd annulus. We are considering the 3rd annulus to be the representative of the whole galaxy in this scenario; more representative than the nucleus or the noisiest annulus, at least. We observe that there is a tendency of larger galaxies to have older age. The slope of the trendline (red color in bottom panel of Fig.  6) is calculated to be 0.59 dex/log($\sigma$). The errorbars associated with each data point on the top panel of Fig.   6 is the mean of the errors for all galaxies within that $\sigma$ bin for a particular annulus. On the other hand, for the bottom panel, the errors for each data point is the error associated with 3rd annulus of each galaxy.
Figure 6: The (log of) mean age is plotted against log of $\sigma$ with errorbars. The top panel shows the mean values for each $\sigma$ bin ($\sigma$ bins are described in Table 2) color coded by annulus. The legend on top shows the middle value for each radial bin. The bottom panel shows the mean age of each galaxy’s 3rd annulus. The legend describes the best fit line with ‘y’ as the mean age and ‘x’ as the log of $\sigma$.
@@
Fig. 7 shows how the [Fe/H] values change with $\sigma$. The top panel shows mean of [Fe/H] values within in $\sigma$ bin for every annulus (color coded). The bottom panel shows the [Fe/H] of the 3rd annulus for each individual galaxy. The negative slope of the binned graph (top panel of Fig. 7) is apparent from the 2nd $\sigma$ bin. When analyzed individually, we find a slightly negative slope overall for the [Fe/H] vs log($\sigma$) graph in bottom panel of Fig. 7, though the two-segment character is still visible in the data, with an inflection point at roughly log $\sigma\approx$ 2.0 or $\sigma\approx$ 100 km s-1.
Figure 7: Mean [Fe/H] is plotted against log $\sigma$ with errorbars. The top panel shows the mean values for each $\sigma$ bin color coded by annulus. The legend on top shows the middle value for each radial bin. The bottom panel shows the mean [Fe/H] of each galaxy for its 3rd annulus. The legend describes the best fit line with ‘y’ as the [Fe/H] and ‘x’ as the log of $\sigma$.
@@ 4.1.3 C, N, Na, and Mg
Having looked into the trends of age and [Fe/H] with local velocity dispersion, we now focus on the elemental abundances of C, N, Na, and Mg for the 2417 MaNGA galaxies.
Figure 8: The mean elemental abundances for C, N, Na, and Mg plotted against log of $\sigma$. The annuli are color coded and values are shown in the legend on top of the figure.
In Fig. 8 the elemental abundances are plotted against log $\sigma$ for each annulus within each $\sigma$ bin. The galaxies are grouped according to their velocity dispersion (at the 3rd annulus) in accordance to Table 2 and the average abundance values are plotted in Fig. 8. The errorbars come from the taking the average of errors of each measurement within a $\sigma$ bin for every annulus. Because the light elements are abundant in comparison with Fe-peak elements, it is clear from this figure that the mean abundance increases with increasing $\sigma$, even if the Fe abundance by itself does not (Fig. 7). Since $\sigma$ correlates with galaxy mass, it is safe to conclude that massive ETGs have significantly higher abundances of C, N, Na, and Mg.
@@
A striking result is obtained, or at least becomes more visible, when we look into elemental abundances at the individual galaxy level instead of grouping them. In Fig. 9, we find the similar trend of increasing abundance with $\sigma$ like Fig. 8. But we find a prominent change of slope at log($\sigma$)=2.0 for all the four elements considered in this work. When we compare the slope values from the trend line equations given in Fig. 9, we see that there is an increase in slope by a factor of 2 for C, 6.2 for N, 2.32 for Na, and 1.6 for Mg for log($\sigma$)$>$2.0. There is a ‘kink’ at log($\sigma$)=2.0 that mirrors the downward inflection for Fe in Fig. 7. These results point to a suspicion that mid-size to large galaxies might follow different evolutionary paths compared to dwarf elliptical galaxies.
Figure 9: The elemental abundances for the 3rd annulus of each galaxy are plotted against $\sigma$. The error bars come from the 50 Monte Carlo realizations as described in the text. The trendlines show the best fit to the data for log($\sigma$)$<$2.0 (in red) and log($\sigma$)$>$2.0 (in cyan). The best fit equations are shown as legend with ‘y’ being the elemental abundance and ‘x’ being the log($\sigma$).
@@ 4.1.4 Binned by [Fe/H]
Apart from investigating the dependence of elemental abundances with $\sigma$, we also looked into how they change with [Fe/H]. Errors in [Fe/H] are included in line fits when using [Fe/H] as an independent variable. Due to the smaller relative errors in velocity dispersion, only the errors in the dependent variables were used when performing regressions using $\sigma$ as the independent variable.
[…]
The consistency of the results up until this point can be checked by comparing Figures 7, 9, and 11. In Fig. 7 the decreasing trend of [Fe/H] with $\sigma$ shows that more massive galaxies have lower [Fe/H] and on the other hand Fig. 9 shows that more massive galaxies have higher light metal abundances. So, it is expected that light metal abundances would decrease with increasing [Fe/H]. This trend is confirmed in Fig. 11. The observed change in slope when light metal abundances are plotted with $\sigma$ vanishes when the same quantities are plotted against [Fe/H], at least for N and Mg. This is an important point to note, as $\sigma$ reflects galaxy structure, while elemental abundances are indicative of chemical evolution processes.
Figure 10: Change of light elements’ abundance plotted against [Fe/H]. The $\sigma$ bins are color coded and shown in the legend on top. Note that the galaxies are grouped by $\sigma$ and [Fe/H] at the 3rd annulus. There are no large galaxies for the highest $\sigma$ and [Fe/H] bin, thus causing absence of data point in the above figure for that bin. Lines of slope $-1$ (black solid line) indicate the behavior if the light element is held at constant abundance while Fe alone increases.
@@
In this work we refrained from stacking galaxies together to explore the results at an individual galaxy level. To this end, it seemed interesting to explore how the light element abundances correlate with each other. In Fig. 12 we plotted the abundances of 3rd annulus of each galaxy against one another. All the elements positively correlate with each other with C perfectly correlating with Na (as evident by the slope of 1.01 in the trendline equation given in Fig. 12). Fig. 12 also corroborates the fact all the elemental abundances follow similar trend when plotted against both $\sigma$ (Fig. 9) and [Fe/H] (Fig. 11).
Figure 12: The elemental abundances are plotted against each other. The abundances are from 3rd annulus of each galaxy. The trendline (in red) equation quantifies their dependence on each other.
@@ 4.2 Gradients in Stellar Population Parameters
In this section we describe our radial gradient results. The radial gradients are calculated in log space by dividing the quantity (log(age), [Fe/H], [C/R], [N/R], [Na/R], and [Mg/R]) by log of the R/Re value. So, the gradient signifies change in any of the quantities (in log scale) for unit change in log(R/Re). The error in radial gradient is calculated from the covariance matrix between the two quantities. Table 4 shows our $\sigma$ binning information along with the mean radial gradients for each bin for age, [Fe/H], and C, N, Na, and Mg abundances.

### v2

@@ 3.2 Composite Model Grid and Parameter Estimation
Naturally, the results of galactic spectral analysis depend on the stellar population model used. The stellar population model utilized for this work is based on Worthey et al. (2014b) (hereafter Worthey model) with the foundational framework established by Worthey (1994). These models use stellar evolutionary isochrones combined with a stellar initial mass function (IMF) to estimate the distribution of stars within the log L versus log Teff diagram. Fluxes are assigned to each star bin, along with empirical estimates of the absorption feature indices.
For any stellar population model three major ingredients are: stellar spectral library, set of isochrones, and initial mass function (IMF). Because synthetic spectra match real ones poorly, spectral libraries usually start with empirical observations with the spectral changes wrought by elemental abundance changes computed synthetically and added later. We adopt this approach. Our empirical library incorporates several stellar spectral libraries: 1. a cross-correlated variant of the Lick spectra (Worthey et al., 2014a), 2. the MILES spectral library (Sánchez-Blázquez et al., 2006), 3. the Indo-US library (Valdes et al., 2004), 4. the IRTF spectral library (Rayner et al., 2003), and 5. the ELODIE spectral library (Prugniel & Soubiran, 2001). These libraries are smoothed or back-corrected to a common 200 km s-1 resolution before indices are measured. Polynomial fitting functions model index behavior as a function of Teff, log($g$), and [Fe/H]. Element sensitivity was added using three spectral synthesis codes as in Lee et al. (2009).
@@
Element sensitivity in $T_{\rm eff}$ was added to the Marigo et al. (2008) isochrones according to the formula of Worthey et al. (2022). We adopted a Kroupa IMF (Kroupa, 2001). To generate an integrated-light index strength, index strengths are calculated for each “star” along the isochrone, and decomposed into an index flux and a continuum flux (weighted by the star’s luminosity at the wavelength of interest and the number of stars in that bin). A running total of the flux and continuum values are then summed over the isochrone. After the summation, the resulting pair of fluxes (continuum and index passbands) is reformulated into a final index value for the single stellar population (SSP). The models used here are not SSPs. They are composite models that incorporate an abundance distribution function (ADF) (Tang et al., 2014; Worthey et al., 2005) that matches the Milky Way and other local galaxies. The models are thus metallicity-composite, but still with a single burst age. In this work we adopt the ‘normal’ width ADF with FWHM = 0.65 dex from Tang et al. (2014). When we quote [Fe/H] results, they refer to the peak metallicity in the asymmetric ADF, not the mass-weighted mean, light-weighted mean, or SSP-equivalent value.
[…]
In a predetermined order, COMPFIT examines various features that are diagnostic of the abundance pattern, weighed by their own sensitivity. With a trial abundance mixture in hand, the program then iterates, improving the age, metallicity, and, separately, chemical element mixture until the model optimally matches each set of galaxy indices. The ”predetermined order” maximizes efficiency compared to relatively blind techniques such as Markov Chain Monte Carlo or Particle Swarm Optimization by starting with unambiguous elements like Na, which affects primarily the Na D index, and proceeds to elements that affect many indices. Grid discretization involving age and metallicity exists, since this portion of the algorithm tests all parts of the parameter space, but the space is parceled into bins. The binning is fine enough so that it is not readily apparent in figures involving age and metallicity. Abundance ratios are linearized, so they do not exhibit discretization.
Once a solution for optimal age and chemical mixture is obtained, the analysis is repeated for copies of the original galaxy index set whose index values have been altered randomly within the observational error envelope. Typically, 50 such Monte Carlo realizations are computed to give output uncertainties. Degeneracies are apparent, especially in the age-metallicity plane, as seen in Figs. 6 and 7. As seen in these figures, the ratios of individual elements generally do not correlate with each other.
Figure 6: Example Monte Carlo parameter inversions for $\sigma\sim 70$ galaxy 11009-12702, innermost annulus. The bold diamond represents the nominal solution, while other points represent parameter solutions from altered versions of 11009-12702.
Figure 7: Example Monte Carlo parameter inversions for $\sigma\sim 200$ galaxy 11965-9101, innermost annulus. Symbols as in Fig. 6.
@@
Here are a few notes on how to interpret the inverted-model results. The final abundance for each element is listed here as the peak (mode) of the ADF, not any kind of SSP-equivalent abundance. Light elements (C, N, O, Si, Na) are assumed to be related to Fe and $Z$ in a single ratio for each solution (that is, no trend of $\alpha$/Fe] with Fe or $Z$ was assumed). Furthermore, due to the subtlety of the spectra features affected by Si and O, inclusion of these elements gives noisy solutions. Therefore, these two elements were tied to Mg, which strongly affects the spectrum. That is, [Mg/Fe]=[Si/Fe]=[O/Fe] always. This matters because isochrone temperatures vary with O and Si abundance, and there is a strong O-age degeneracy (Worthey et al., 2022). Elements not mentioned, such as Ni, Cr, Pd, or U, were similarly tied in lockstep to Fe. We refer to such unnamed elements by the symbol ‘R.’ Helium abundance has a non-adjustable trajectory with $Z$ determined by the isochrone set. This is out of pragmatic necessity, since He cannot be measured in the galaxies studied here, but should be mentioned since it puts an irreducible systematic error into the results.
@@ 4 Results
This section presents the results for stellar age, [Fe/H], and the abundances of C, N, Na, and Mg, along with their radial gradients. For the purposes of this study, we refer to the single-burst-equivalent mean age, [Fe/H], [C/R], [N/R], [Na/R], and [Mg/R] collectively as stellar population parameters. In $\S$4.1, we show the variation of these parameters as a function of R/Re, out to R/Re = 0.7. Average values of the parameters, binned by velocity dispersion ($\sigma$), are reported in Table 2, while Table 3 presents the same parameters binned by [Fe/H]. In both cases, the tabulated values correspond to the mean measurements from the 3rd annulus of galaxies within each bin. Section 4.2 presents the results on the radial gradients of the stellar population parameters.
@@ 4.1 Stellar Population Parameters
Stellar population parameters were derived using compfit, as described in $\S$3.2. Uncertainties were estimated by performing 50 Monte Carlo (MC) realizations per annulus for each galaxy. In each realization, the measured, corrected index strengths were perturbed within a Gaussian envelope of 0.7 SD (Standard Deviation). The 0.7 factor was chosen to suppress outlying solutions and improve error robustness in some cases. The standard deviation across the 50 realizations was adopted as the parameter uncertainty for each annulus. While [C/R], [N/R], [Na/R], and [Mg/R] solutions are treated as linear and unbounded, age and [Fe/H] can fall near the model grid edges—specifically, [Fe/H] near 0.4 and age near 16 Gyr. For solutions at or near these boundaries, MC-derived errors were linearly inflated, up to a factor of two, to account for the absence of solutions beyond the grid limits.
@@ 4.1.1 Age
Since mass is proportional to $\sigma^{2}$ but scales linearly with galaxy size (Bender et al., 1992), $\sigma$ serves as an imperfect yet sufficient proxy for galaxy mass. Fig. 8 shows the variation of log(mean age [Gyr]) with $\sigma$. Ages derived in this work are, roughly, $U$-band-weighted mean ages. In the top panel, the mean ages (represented by the center of the rectangles) across $\sigma$ bins (color-coded by annulus) are plotted against the log of $\sigma$, where the $\sigma$ values represent bin centers as defined in Table 2. It is to be noted that the height of each rectangular box in top panel of Fig. 8 corresponds to error for a given annulus at a given $\sigma$ whereas the width of the box, although arbitrarily chosen, tracks radius. The bottom panel displays the age at the 3rd annulus for each galaxy, which we adopt as a representative value, more so than the central region or outermost, noisier annuli. A positive trend is apparent, indicating that galaxies with higher $\sigma$ tend to be older. The slope of the best-fit line (shown in red in the bottom panel) is 0.53 dex per log $\sigma$. The error bars in the top panel reflect the average parameter uncertainty across all galaxies within a $\sigma$ bin for each annulus, while in the bottom panel they correspond to the uncertainty in age at the 3rd annulus of each individual galaxy.
[…]
The legend on top lists the middle value for each radial bin. The bottom panel shows the mean age (with errorbars) of each galaxy’s 3rd annulus. The legend describes the best fit line with ‘y’ as the mean age and ‘x’ as the log of $\sigma$.
@@
[…]
The legend on top lists the middle value for each radial bin. The bottom panel shows the mean [Fe/H] (with errorbars) of each galaxy for its 3rd annulus. The legend describes the best fit line with ‘y’ as the [Fe/H] and ‘x’ as the log of $\sigma$.
@@ 4.1.3 C, N, Na, and Mg
Having examined the trends of age and [Fe/H] with local velocity dispersion, we now turn to the elemental abundances of C, N, Na, and Mg in our sample of 2968 MaNGA galaxies.
Figure 10: The mean elemental abundances for C, N, Na, and Mg plotted against log of $\sigma$. As in Figs. 8 and 9, the center of the rectangular box corresponds to mean value and the error is represented by the height of the box.
The annuli are color coded and annulus midpoints are listed in the legend on top of the figure.
The galaxies are grouped according to their $\sigma$ (at the 3rd annulus) according to Table 2 and the average abundance values in each $\sigma$ bin are plotted in Fig. 10 (color-coded by annulus). Note that the center of the rectangular box represents the average abundance value while the height represents the corresponding error. The errors are obtained by taking the average of the error of each measurement within a $\sigma$ bin for every annulus. Given the greater abundance of light elements relative to Fe-peak elements, this figure clearly illustrates that the mean abundance rises with increasing $\sigma$, even though the Fe abundance alone does not (Fig. 9). As $\sigma$ is correlated with galaxy mass, it can be safely concluded that massive ETGs exhibit significantly elevated abundances of C, N, Na, and Mg compared to their less massive counterparts.
@@
A notable finding emerges when examining elemental abundances at the individual galaxy level rather than in groups. As depicted in Fig. 11, a consistent trend of increasing abundance with $\sigma$ is observed, mirroring the pattern in Fig. 10. However, a distinct change in slope is evident at log($\sigma$)=2.0 for all four elements analyzed in this study. Analysis of the trend line equations in Fig. 11 reveals that the slope increases by a factor of 2.6 for C, 3.26 for N, 2.26 for Na, and 1.3 for Mg for log $\sigma>$2.0 compared to log $\sigma<$2.0. This transition at log($\sigma$)=2.0 corresponds to the downward inflection in Fe abundance, shown as a ‘kink’ in Fig. 9. These observations suggest that mid-sized to large galaxies may follow distinct evolutionary pathways compared to dwarf elliptical galaxies.
Figure 11: The elemental abundances for the 3rd annulus of each galaxy are plotted against $\sigma$. The error bars come from the 50 Monte Carlo realizations as described in the text. The trendlines show the best fit to the data for log($\sigma$)<<2.0 (in red) and log($\sigma$)>>2.0 (in cyan). The best fit equations are shown as legend with ‘y’ being the elemental abundance and ‘x’ being the log($\sigma$).
@@ 4.1.4 Binned by [Fe/H]
In addition to examining the relationship between elemental abundances and $\sigma$, we investigated their variation with respect to [Fe/H]. As [Fe/H] is derived from the compfit inversion program, we incorporate the associated uncertainties in [Fe/H] when charting the variation of other parameters with [Fe/H].
Fig. 12 presents a plot of elemental abundances as a function of [Fe/H]. In this figure, we plotted the average abundance value across each [Fe/H] bin color-coded by $\sigma$. The error bars represent the mean of uncertainties within each [Fe/H] bin across all $\sigma$ bins, reflecting a typical single-galaxy error. A general decreasing trend in light element abundances is observed with increasing [Fe/H]. Trend lines with a slope of 1-1 (solid black lines in Fig. 12) illustrate the expected behavior if light element abundances remain constant while Fe increases, as would occur if Type Ia supernovae were the sole contributors. The observed shallower decline in light element abundances suggests ongoing enrichment of these elements at higher [Fe/H]. This trend is further quantified in Fig. 13, where the elemental abundances at the 3rd annulus of each galaxy are plotted against [Fe/H]. Linear trend lines fitted to the data exhibit slightly negative slopes ($\sim$ -0.1 dex per decade) for all elements (see Fig. 13). A slope of zero would indicate lockstep enrichment with Fe, whereas a slope of 1-1 would suggest no enrichment. The observed slopes are significantly closer to the lockstep scenario. Romano et al. (2010) documents the behavior of various elements with respect to [Fe/H]. Within the [Fe/H] regime of -0.6<<[Fe/H]<<0.4, our findings for C, N, Na, and Mg very closely resonates with the findings of Romano et al. (2010).
The consistency of the findings can be verified by comparing Figures 9, 11, and 13. Fig. 9 illustrates a decreasing trend in [Fe/H] with increasing $\sigma$, indicating that more massive galaxies exhibit lower [Fe/H]. Conversely, Fig. 11 demonstrates that more massive galaxies possess higher abundances of light metals. Consequently, it is anticipated that light metal abundances would decrease with increasing [Fe/H], a trend confirmed in Fig. 13. Notably, the change in slope observed when plotting light metal abundances against $\sigma$ disappears when these abundances are plotted against [Fe/H], at least for N and Mg. This distinction is significant, as $\sigma$ reflects galaxy structure, whereas elemental abundances are indicative of chemical evolution processes.
Figure 12: Change of light elements’ abundance plotted against [Fe/H]. The $\sigma$ bins are color coded and shown in the legend on top. Note that the galaxies are grouped by $\sigma$ and [Fe/H] at the 3rd annulus. There are no large galaxies for the highest $\sigma$ and [Fe/H] bin, thus causing absence of data point in the above figure for that bin. Lines of slope 1-1 (black solid line) indicate the behavior if the light element is held at constant abundance while Fe alone increases.
@@ 4.2 Gradients in Stellar Population Parameters
This section presents the results of our radial gradient analysis. Radial gradients are computed in logarithmic space by dividing the stellar population parameters (age, [Fe/H], [C/R], [N/R], [Na/R], and [Mg/R]) by the logarithm of the normalized radius, log(R/Re). Thus, the gradient represents the change in a given parameter (on a logarithmic scale) per unit change in log(R/Re). The associated uncertainties in the radial gradients are derived from the covariance matrix of the paired quantities. Table 4 provides the details of $\sigma$ binning, alongside the mean radial gradients for each bin for age, [Fe/H], and the abundances of C, N, Na, and Mg.